OpenStudy (anonymous):

For DrPepper: I tried to anwer it but it seems to be hung up. It is for the problem dealing withthe equation of the line y=mx+b. I came up with 22/3. They want the answer to be "slope intercept form"

OpenStudy (anonymous):

sorry, I mean 17/3

OpenStudy (bahrom7893):

post the original question

OpenStudy (bahrom7893):

ok. m=5/6, (2,-4) and the answer needs to be in slope-intercept form

OpenStudy (bahrom7893):

Okay so slope intercept is y = Mx+b, but ignore that for now. I'm just gonna use this formula: y-y1 = M(x-x1) y - (-4) = (5/6)(x-2)

OpenStudy (bahrom7893):

y + 4 = (5/6)x - (5/3)

OpenStudy (bahrom7893):

y = (5/6)x - (5/3) - 4 y = (5/6)x - 29/6

OpenStudy (bahrom7893):

Why is everyone on this two year old question again?

OpenStudy (anonymous):

because time travel

OpenStudy (bahrom7893):

oh that's what u posted mod lol

OpenStudy (anonymous):

trying to find oldest qustion

OpenStudy (bahrom7893):

I doubt you will... i tried going back a while ago, couldn't. But i got the start date to be Feb 20, 2011